# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation
1:Input: MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$ with decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$.
@@
3:  Query $\mathbb{A}$ and receive $u^{n}\in\Delta(\mathcal{B})$.
4:  Execute $\pi^{n}:=\pi_{u^{n}}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
@@
Algorithm 3 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 24). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ whose mixing weight $u^{n}$ is obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Same as Algorithm 1, Algorithm 3 also requires $\mathbb{A}$ to have decision space $\Delta(\mathcal{B})$ and benchmark set $\left\{e_{\pi}:\pi\in\mathcal{B}\right\}$. It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(\pi)$ (Eq. (5)). At the end of round $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{H}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 3:
@@
Theorem 27 shows that the interactive IL Algorithm 3 matches the trajectory-wise sample complexity of behavior cloning in [foster2024behavior]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [foster2024behavior, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^3] For [foster2024behavior, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\log|\mathcal{B}_{h}|$ to be significantly smaller $R\log|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu<R/H<1$ in the practically-popular parameter sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
For the proof of Theorem 27, we introduce a new notion of decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependence between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, it is compatible with non-Markovian first-step mixing of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest.
@@ E.3 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi}(s_{1:H},a_{1:H}),\mathbb{P}^{\pi^{\mathrm{E}}}(s_{1:H},a_{1:H})\right)$ [foster2024behavior], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]$ that decouples states and actions is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemma shows that such relationship holds for both Markovian policies and their first-step mixings.
@@
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi$ be an Markovian policy. Then we have
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
To prove these two lemmas, we first prove a special case, i.e. Lemma 30 with first-step mixing of deterministic policies. To facilitate the proofs, we introduce the following additional notations:
  • Let $\mathcal{B}^{\mathrm{Det}}$ denote the set of all deterministic, Markovian policies. We will use $\nu,\nu^{\prime}$ to denote members of $\mathcal{B}^{\mathrm{Det}}$ and $\nu_{h}(s)$ to denote the action taken by $\nu$ at $(s,h)$ when it is clear from the context.
@@
Lemma 30.
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixing of deterministic Markovian policies (elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
The key idea in the following proof is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (9)).
@@
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 35, yielding:
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$
@@
$\frac{1}{2}\parallel\pi_{u}(s_{1:H})-\pi^{\mathrm{E}}(s_{1:H})\parallel_{1}=1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})),$
@@
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H})).$ (7)
@@
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$ (8)
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
@@
Given deterministic Markovian policies $\nu$, $\nu^{\prime}$, and $\pi^{\mathrm{E}}$, the following holds
@@
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (12)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}P_{0}(\mathbf{E^{\prime}})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
@@
E.3.1 Proof of Lemma 28
Proof.
The key to this proof is showing that any Markovian policy $\pi$ is equivalent, in terms of action distribution on any state sequence, to a first-step mixing of a set of deterministic Markovian policies. This leads to equivalence on trajectory distribution and decoupled Hellinger distance. To clarify further, we present the following claim, which allows us to apply guarantees for mixtures of deterministic policies in Lemma 30.
Claim 32.
For a Markovian policy $\pi$, there exists a first-step mixing of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(s_{1:H})=\pi_{u}(s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
Given an MDP with finite state space size $S$ and action space size $A$, the set of all deterministic, Markovian policies, denoted by $\mathcal{B}^{\mathrm{Det}}$, contains $A^{SH}$ deterministic policies, which can be indexed by a tuple of actions $(a_{h,s})_{h\in[H],\,s\in\mathcal{S}}$.
To construct policy $\pi_{u}$, we will set the weight vector $u$ such that its weight on policy $\nu$ indexed by $(a_{h,s})_{h\in[H],s\in{\mathcal{S}}}$ as:
$u(\nu)=\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a_{h,s}|s)$ (19)
It can be easily verified by that $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)=1$.
$\pi_{u}(a_{1:H}\parallel s_{1:H})=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\prod_{h=1}^{H}\nu(a_{h}|s_{h})$ (20)
$=$ $\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})=\pi(a_{1:H}\parallel s_{1:H}).$
Since this holds for any action sequence $a_{1:H}\in\mathcal{A}^{H}$, we derive the first part of Claim 32 that $\pi(s_{1:H})=\pi_{u}(s_{1:H})$. The second item follows from the first item in combination with Lemma 22.
We conclude that for the $\pi_{u}$ in the statement of the claim,
$\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]=\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
Finally, the proof follows by applying Lemma 30 to $\pi_{u}$. ∎
E.3.2 Proof of Lemma 29
Proof.
By Claim 32, any Markovian policy can be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, then any first-step mixing of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof follows by applying Lemma 30. ∎
E.4 New Guarantees for DAgger Variant with Trajectory-wise Annotation
Recall that we have defined decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right].$
In the following, we first demonstrate that the performance difference between expert and the the uniform first-step mixing of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper-bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 3 in Theorem 27.

### v3

@@ E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation Sample Complexity
1:Input: MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, stationary policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$ with decision space $\Pi_{\mathcal{B}}$.
@@
3:  Query $\mathbb{A}$ and receive $\pi^{n}\in\Pi_{\mathcal{B}}$.
4:  Roll out $\pi^{n}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
@@
Algorithm 3 uses first-step mixing policies $\pi_{u}\in\Pi_{\mathcal{B}}$ (recall Definition 28). At round $n$, it rolls out $\pi^{n}=\pi_{u^{n}}$ obtained from an online learning oracle $\mathbb{A}$ and samples a full state sequence $s^{n}_{1:H}$. Similar to Algorithm 1, Algorithm 3 also requires $\mathbb{A}$ to have decision space $\Pi_{\mathcal{B}}$ (cf. $\bar{\Pi}_{\mathcal{B}}$ in Algorithm 1). It then requests expert’s trajectory-wise annotation $a^{*,n}_{1:H}$ and updates $\mathbb{A}$ by $\ell^{n}(\pi)$ (Eq. (13)). At the end of iteration $N$, the uniform first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ is returned, which is equivalent to returning $\pi_{\hat{u}}$, where $\hat{u}:=\frac{1}{N}\sum_{n=1}^{N}u^{n}$. We provide the following performance guarantee of Algorithm 3:
@@
Theorem 31 shows that in the deterministic realizable setting, the interactive IL Algorithm 3 has a trajectory-wise sample complexity matching that of behavior cloning [17]. In contrast, prior state-of-the-art analysis of interactive IL algorithms [17, Appendix C.2] gives sample complexity results that are in general worse than behavior cloning.  [^6] For [17, Appendix C.2]’s sample complexity to improve over behavior cloning, we need $\mu H\max_{h\in[H]}\log|\mathcal{B}_{h}|$ to be significantly smaller $R\log|\mathcal{B}|$ (where $\mathcal{B}_{h}$ is the projection of $\mathcal{B}$ onto step $h$). This may require the strong condition that $\mu\ll R/H\leq 1$ in the more practical parameter-sharing settings ($|\mathcal{B}_{h}|=|\mathcal{B}|$).
For the proof of Theorem 31, we introduce a new notion called decoupled Hellinger estimation error:
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}:=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].$
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}$ decouples the dependence between the state sequence and the distribution of action sequence induced by the learner. Perhaps surprisingly, as we show below, it is compatible with non-Markovian first-step mixing of policies, while still being well-behaved enough to be translated to a policy suboptimality guarantee, which could be of independent interest.
@@ E.3 Decoupling State and Action Sequences by Decoupled Hellinger Distance
In this section, we demonstrate that similar to $D^{2}_{\text{H}}\left(\mathbb{P}^{\pi},\mathbb{P}^{\pi^{\mathrm{E}}}\right)$ [17], the decoupled Hellinger distance $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right]$ is also proportionally lower bounded by a constant factor of $\rho(\pi\parallel\pi^{\mathrm{E}})$. The following two lemmas show that such relationship holds for both Markovian policies and their first-step mixings.
@@
Let $\pi^{\mathrm{E}}$ be a deterministic policy, and let $\pi$ be a Markovian policy. Then we have
$\frac{1}{2}\cdot\rho(\pi\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].$
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].$
The cornerstone of the above two lemmas is the following special case about first-step mixing of deterministic policies. Given an MDP with finite state space size $S$ and action space size $A$, we denote the set of all deterministic, Markovian policies, as $\mathcal{B}^{\mathrm{Det}}$. $\mathcal{B}^{\mathrm{Det}}$ contains $A^{SH}$ deterministic policies, which can be indexed by a tuple of actions $(a_{h,s})_{h\in[H],\,s\in\mathcal{S}}$.
Lemma 34.
Let $\pi^{\mathrm{E}}$ be a deterministic Markovian policy, and let $\pi_{u}$ be a first-step mixing of deterministic Markovian policies (i.e., elements of $\mathcal{B}^{\mathrm{Det}}$). Then we have that
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].$
We now quickly conclude Lemmas 32 and 33 using Lemma 34 in the next subsection, then come back to the proof of Lemma 34 in the subsection after.
E.3.1 Proofs of Lemmas 32 and 33
Proof of Lemma 32.
We show the following simple claim: any Markovian policy $\pi$ is equivalent to a first-step mixing of a set of deterministic Markovian policies. This allows us to apply guarantees for mixtures of deterministic Markovian policies in Lemma 34 to conclude the proof. ∎
Claim 35.
For a Markovian policy $\pi=(\pi_{1},\ldots,\pi_{H})$, there exists a first-step mixing of deterministic policy $\pi_{u}$ such that for any $s_{1:H}\in{\mathcal{S}}^{H}$, 1. $\pi(\cdot\|s_{1:H})=\pi_{u}(\cdot\|s_{1:H})$, and 2. $\mathbb{P}^{\pi}(s_{1:H})=\mathbb{P}^{\pi_{u}}(s_{1:H})$.
Proof of Claim 35.
To construct policy $\pi_{u}$, we will set the weight vector $u$ (over $\mathcal{B}^{\mathrm{Det}}$) such that its weight on policy $\nu$ indexed by $(a_{h,s})_{h\in[H],s\in{\mathcal{S}}}$ as:
$u(\nu)=\prod_{h=1}^{H}\prod_{s\in{\mathcal{S}}}\pi_{h}(a_{h,s}|s)$ (14)
It can be easily verified that $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)=1$.
$\pi_{u}(a_{1:H}\parallel s_{1:H})=$ $\sum_{\nu\in\mathcal{B}^{\mathrm{Det}}}u(\nu)\prod_{h=1}^{H}\nu_{h}(a_{h}|s_{h})$ (15)
$=$ $\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})=\pi(a_{1:H}\parallel s_{1:H}).$
Since this holds for any action sequence $a_{1:H}\in\mathcal{A}^{H}$, we derive the first part of Claim 35 that $\pi(\cdot\|s_{1:H})=\pi_{u}(\cdot\|s_{1:H})$. The second item follows from the first item combined with Lemma 15. ∎
Proof of Lemma 33.
By Claim 35, any Markovian policy can be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$, thus, any first-step mixing of Markovian policies $\pi_{u}$ can also be viewed as a first-step mixing of $A^{SH}$ deterministic policies from $\mathcal{B}^{\mathrm{Det}}$. The proof again follows by applying Lemma 34. ∎
E.3.2 Proof of Lemma 34
To facilitate the proof of Lemma 34, we introduce the following additional notations:
  • Recall that $\mathcal{B}^{\mathrm{Det}}$ is the set of all deterministic, Markovian policies. We will use $\nu,\nu^{\prime}$ to denote members of $\mathcal{B}^{\mathrm{Det}}$ and $\nu_{h}(s)$ to denote the action $\nu$ takes at state $s$ at step $h$ when it is clear from the context.
@@ E.3.2 Proof of Lemma 34
Our key idea is to lower bound $\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right],$ which reflects the asymmetric roles of the two appearances of $\pi_{u}$’s, using a symmetric formulation via function $F$ (as shown in (19) below).
@@
Next, we turn to analyzing $D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))$. Since the deterministic expert induces a delta mass distribution over actions, we apply the elementary fact about the Hellinger distance with delta mass distribution stated in Lemma 38, yielding:
$\frac{1}{2}\parallel\pi_{u}(\cdot\|s_{1:H})-\pi^{\mathrm{E}}(\cdot\|s_{1:H})\parallel_{1}\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H})).$
@@
$\frac{1}{2}\parallel\pi_{u}(\cdot\|s_{1:H})-\pi^{\mathrm{E}}(\cdot\|s_{1:H})\parallel_{1}=1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})),$
@@
$1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H}))\leq D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H})).$ (17)
@@
$\sum_{s_{1:H}}\mathbb{P}^{\pi_{u}}(s_{1:H})(1-u(\mathcal{B}^{\mathrm{E}}(s_{1:H})))\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].$ (18)
@@
$\frac{1}{2}\cdot\rho(\pi_{u}\parallel\pi^{\mathrm{E}})\leq(*)\leq\mathbb{E}^{\pi_{u}}\left[D^{2}_{\text{H}}(\pi_{u}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right].\qed$
@@
Given deterministic Markovian policies $\nu$, $\nu^{\prime}$, and $\pi^{\mathrm{E}}$, the following holds:
@@
$\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu}(s_{1:h})=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\rho(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu_{h}(s_{h}))$ (22)
$=$ $\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\rho(s_{1})\prod_{h^{\prime}=1}^{h-1}P_{h}(s_{h+1}|s_{h},\nu^{\prime}_{h}(s_{h}))=\sum_{s_{1:h}\in\tilde{\mathcal{X}}_{i}^{h}}\mathbb{P}^{\nu^{\prime}}(s_{1:h}).$
@@
E.4 Proof of Theorem 31
We first demonstrate that the performance difference between expert and the the uniform first-step mixing of any Markovian policy sequence $\{\pi^{n}\}_{n=1}^{N}$ is upper bounded by $2R\operatorname{OnEst}^{\mathrm{Traj}}_{N}/N$, and then show the trajectory-wise sample complexity of Algorithm 3 in Theorem 31.